The vector aā + Bb where a = 5, B =-2, à = (9, -3), b =(-1, 2) is parallel to
Answer #1

Hey there stay home and stay safe

\small \alpha=\frac{1}{3},\beta=-2,\vec a=(9,-3),\vec b=(-1,2)\\ \alpha \vec a+\beta \vec b\\ =\frac{1}{3}(9,-3)-2(-1,2)\\ =(3,-1)+(2,-4)\\ =(5,-5)\\ =5(1,-1)\\ \alpha \vec a+\beta \vec b$ is parallel to $(1,-1)

And please leave a like. Thank you.